tana233:

What is the sign of the product (−2) × (4) × (−6) × (−3)? Positive, because the products (−2) × (4) and (−6) × (−3) are negative and the product of two negative numbers is positive Positive, because the products (−2) × (4) and (−6) × (−3) are positive and the product of two positive numbers is positive Negative, because the product (−2) × (4) is positive and the product (−6) × (−3) is negative and the product of a positive number and a negative number is negative Negative, because the product (−2) × (4) is negative and the product (−6) × (−3) is positive and the product of a negative number and a positive number is negative

Eiwoh2:

All you have to do in this one is firstly multiply both (-2)x(4) and (-6)x(-3). Multiplying (-2)x(4) gets us 18, a positive number. And multiplying (-6)x(-3) gets us -8, a negative number. Now if we multiply (−2) × (4) × (−6) × (−3) all together, it gives us a negative number: -144. Meaning, it would be negative. But which of the last two options would you guess is the answer based on this work?

TheSmartOne:

Remember When you multiply a positive number with a positive number, you get a positive number When you multiply a negative number with a negative number, you get a positive number When multiply a positive number with a negative number, you get a negative number And vice versa. i.e when you multiply a negative number with a positive number, you get a negative number Summarized (+) * (+) = (+) (-) * (-) = (+) (+) * (-) = (-) (-) * (+) = (-)